Riddle Industries has the following parameters related to its stock and firm.

Beta: 1.5
Recent Dividend 0.70 dollars
Dividend Growth Rate 4.5 percent
Expected return on market 12.5 percent
Treasury Bills Yield 4.3 percent
Most recent stock price 64.00 dollars

What is the cost of equity using DDM? What is the cost of equity using SML?

A. 5.64 percent, 16.60 percent

B. 5.44 percent, 12.30 percent

C. 4.5 percent, 12.50 percent

D. 5.72 percent, 16.60 percent
Answer #1

A. 5.64 percent, 16.60 percent

Cost of equity using DDM = (D0*(1+g)/P0)+g Where,
= (0.70*(1+0.045)/64.00)+0.045 D0 = Recent dividend = $       0.70
= 5.64% g = Dividend growth rate = 4.50%
P0 = Most recent stock price = $    64.00
Cost of equity using SML = Risk free rate+beta*(Market return - Risk free rate)
= 4.3%+1.5*(12.5%-4.3%)
= 16.60%